OpenStudy (anonymous):

In triangle DEF, angle D = 44 degrees, angle E = 61 degrees, and segment EF = 20 inches. What is the length of segment DE?

OpenStudy (anonymous):

So you can add the two angles together then subtract by 180 first

OpenStudy (anonymous):

angle f = 75

OpenStudy (anonymous):

Then do law of sines.

OpenStudy (anonymous):

Laws of sines. You have the angle D (opposite to EF), just discover the angle F (opposite to DE). 44 + 61 + angle F = 180 angle F= 180-44-61 = 75 Now, laws of sines: EF/sin44º = ED/sin75º

OpenStudy (anonymous):

so sin44/20=sin75/DE

OpenStudy (anonymous):

sin44/20 = . 035, sin75/de = .035

OpenStudy (anonymous):

Just solve for DE.

OpenStudy (anonymous):

DE = sin75 x sin44/20?

OpenStudy (anonymous):

No, DE = 20*sin75/sin44

OpenStudy (anonymous):

27.8

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

A cruise ship travels 300 miles east, then turns 30 degrees north, and travels another 175 miles. What is the total distance from its starting point?

OpenStudy (anonymous):

Laws of cosines. You have 2 sides and the angle between this sides. Just use the formula that I said to you yesterday: c² = a² + b² - 2abcos theta c is the side that you want to discover opposite to the angle theta.

OpenStudy (anonymous):

172 miles

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

In triangle XYZ, XY = 15, YZ = 21, and XZ = 27. What is the measure of angle Z to the nearest degree?

OpenStudy (anonymous):

xy is opposite of angle Z

OpenStudy (anonymous):

XZ is hypotenuse, YZ is adjacent

OpenStudy (anonymous):

the triangle is right ?

OpenStudy (anonymous):

i don't know, there's no picture or diagram

OpenStudy (anonymous):

So you can't that there's a hypotenuse rs.

OpenStudy (anonymous):

cant say*

OpenStudy (anonymous):

oh, okay. i just assumed the hypotenuse was XZ because it's the longest segment

OpenStudy (anonymous):

you have to use the other laws of cosine formula. Cos theta = (a² + b² - c²) / ( 2*a*b)

OpenStudy (anonymous):

c is the side opposite to the angle that you want.

OpenStudy (anonymous):

so XY is my c

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i got 89.9 but that's not an option

OpenStudy (anonymous):

are you sure? cos theta = (27² + 21² - 15² )/2*27*21 = 0.83

OpenStudy (anonymous):

do arccos with 0.83

OpenStudy (anonymous):

33.9 which rounds up to 34

OpenStudy (anonymous):

there's an option?

OpenStudy (anonymous):

yep! 34 is an option

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

verify the identity: http://puu.sh/3e9qD.png

OpenStudy (anonymous):

It's to say that it's true or false?

OpenStudy (anonymous):

Or re-do the identity ?

OpenStudy (anonymous):

i have to show the steps of verifying that it is true or false

OpenStudy (anonymous):

Ok, okay,

OpenStudy (anonymous):

a minute

OpenStudy (anonymous):

I got it, i will write in latex.

OpenStudy (anonymous):

\[\frac{ \sec x }{ \csc x - \cot x } - \frac{ \sec x }{ \csc x + \cot x } = \frac{ \sec x (\csc x + \cot x) - \sec x (\csc x - \cot x) }{ \csc²x - \cot²x }\] \[\csc² x - \cot ²x = 1\] So, we will have, \[\sec x(\csc x + \cot x) - \sec x (\csc x - \cot x) =\] \[\sec x \csc x + \sec x \cot x - \sec x \csc x + \sec x \cot x\]\[2\sec x \cot x = 2 \frac{ 1 }{ \cos x } * \frac{ 1 }{ \tan x } = 2 \frac{ 1 }{ \cos x } * \frac{ \cos x }{ sen x } = \]\[2*\frac{ 1 }{ sen x } = 2\csc x\]

OpenStudy (anonymous):

If you don't get it, i will try to explain better.

OpenStudy (anonymous):

i get it. thank you so much.

OpenStudy (anonymous):

Just to make it complete, I said that csc²x - cot²x = 1 because: \[\csc² x = \frac{ 1 }{ sen² x } \iff \cot² x = \frac{ \cos² x }{ sen² x }\]\[\csc² x - \cot² x = \frac{ 1 }{ sen² x } - \frac{ \cos² x }{ sen² x } = \frac{ 1- \cos² x }{ sen x }\]But you now that \[sen² x + \cos² x = 1\]So,\[sen²x = 1 - \cos²x\]To conclude,\[\csc²x - \cot²x = \frac{ 1 - \cos²x }{ sen²x } = \frac{ sen²x }{ sen²x } = 1\]